OpenStudy (alivejeremy):

Andy estimated that he would need 71 feet of lumber for a tree house project. He later found that the actual amount of lumber needed was 60 feet. What was the percent error of Andy's estimation? 15.49% 18.33% 11% 84.51%

OpenStudy (alivejeremy):

Need answer fast :D

OpenStudy (minimato):

(71- 60)/71 = 11/71 = 0.1549 = 15.49%

OpenStudy (alivejeremy):

No it's 18.33%

OpenStudy (alivejeremy):

;)

OpenStudy (minimato):

Oh well, sorry about that

OpenStudy (minimato):

I don't do well when rushed ;)